The data set given below consists of six pairs of (x, y); (10, 70); (12, 65); (2, 96); (0, 94); (8, 75); (5, 82) I Based on a plot, determine whether the relationship between x and y is linear or not. a. b. Find the value of r to analyze the strength of the relationship.


1
Expert's answer
2021-06-10T08:39:54-0400

a.


"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c:c}
 & x & y & xy & x^2 & y^2 \\ \hline
 & 10 & 70 & 700 & 100 & 4900 \\
 \hdashline
 & 12 & 65 & 780 & 144 & 4225 \\
 \hdashline
 & 2 & 96 & 192 & 4 & 9216 \\
 \hdashline
 & 0 & 94 & 0 & 0 & 8836 \\
 \hdashline
& 8 & 75 & 600 & 64 & 5625 \\
 \hdashline
& 5 & 82 & 410 & 25 & 6724 \\
 \hdashline
Sum=& 37 & 482 & 2682 & 337 & 39526 \\
 \hdashline
\end{array}"


"\bar{x}=\dfrac{\displaystyle\sum_{i=1}^nx_i}{n}=\dfrac{37}{6}"

"\bar{y}=\dfrac{\displaystyle\sum_{i=1}^ny_i}{n}=\dfrac{482}{6}"

"SS_{xx}=\displaystyle\sum_{i=1}^nx_i^2-\dfrac{1}{n}\big(\displaystyle\sum_{i=1}^nx_i\big)^2=\dfrac{653}{6}"

"SS_{yy}=\displaystyle\sum_{i=1}^ny_i^2-\dfrac{1}{n}\big(\displaystyle\sum_{i=1}^ny_i\big)^2=\dfrac{4832}{6}"

"SS_{xy}=\displaystyle\sum_{i=1}^ny_i^2-\dfrac{1}{n}\big(\displaystyle\sum_{i=1}^nx_i\big)\big(\displaystyle\sum_{i=1}^ny_i\big)=-\dfrac{1742}{6}"

"m=\dfrac{SS_{xy}}{SS_{xx}}=-\dfrac{1742}{653}\approx-2.6677"

"n=\bar{y}-m\bar{x}=\dfrac{482}{6}+\dfrac{37}{6}\cdot\dfrac{1742}{653}\approx96.7841"

"Y=96.7841-2.6667X"


The relationship between x and y is linear. 


b.


"r=\dfrac{SS_{xy}}{\sqrt{SS_{xx}}\sqrt{SS_{yy}}}=\dfrac{-\dfrac{1742}{6}}{\sqrt{\dfrac{653}{6}}\sqrt{\dfrac{4832}{6}}}\approx-0.9807"

"r^2\approx0.9617"

"0.7<r\leq 1"

Strong negative correlation.